OpenStudy (anonymous):

Can anyone please help me with this? What is the quotient of t+3/2t ÷ 2t+6/t-4?

OpenStudy (anonymous):

t-3 (t+4/3) =2t +3 t-3t-4=2t+3 -4t=7 t=-7/4